News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 12 November 1999 9.00am EST

During the last 48 hours, Chandra completed observations of Capella, and Cygnus X-2, the observatory passed through the radiation belts with associated ACIS CTI measurements, and a calibration observation of HZ-43 was started.

Real-time commanding was completed as planned at 1am on 11/10 to remove the ACIS engineering patch used to investigate a fix for the FEP-0 parity error problem. The patch was installed for approximately 24 hours and an initial look at the data suggests that the problem did not reoccur during the test period. A further test is expected.

During the next 24 hours, calibration observations will continue of HZ-43 and the Coma cluster with HRC-S/LETG and HRC-I and, and GTO observation of the Hubble Deep Field will be conducted with ACIS-I.

Plans were finalized for the Leonids meteor show transition on 11/17, and incorporated into the next mission plan starting 8am 11/14.

The spacecraft subsystems continued to operate nominally.





Return to Status Reports